And can you show me how to calculate the P-value please
Hw Score: 40%, 10of25pes 11.3.8-T H: A white candidale is more likely to pass the test than a minority candidale
Answer #1

Ans:

Observed(fo)
Passed failed Total
White 15 16 31
Minority 9 26 35
Total 24 42 66
Expected(fe)
Passed failed Total
White 11.27 19.73 31
Minority 12.73 22.27 35
Total 24 42 66
Chi square=(fo-fe)^2/fe
Passed failed Total
White 1.232 0.704 1.937
Minority 1.092 0.624 1.715
Total 2.324 1.328 3.652

Test statistic:

Chi square=3.652

df=(2-1)*(2-1)=1

p-value=CHIDIST(3.652,1)=0.0560